# ClassContrast: Bridging the Spatial and Contextual Gaps for Node Representations

### v1

@@ Abstract
Graph Neural Networks (GNNs) have revolutionized the domain of graph representation learning by utilizing neighborhood aggregation schemes in many popular architectures, such as message passing graph neural networks (MPGNNs). This scheme involves iteratively calculating a node’s representation vector by aggregating and transforming the representation vectors of its adjacent nodes. Despite their effectiveness, MPGNNs face significant issues, such as oversquashing, oversmoothing, and underreaching, which hamper their effectiveness. Additionally, the reliance of MPGNNs on the homophily assumption, where edges typically connect nodes with similar labels and features, limits their performance in heterophilic contexts, where connected nodes often have significant differences. This necessitates the development of models that can operate effectively in both homophilic and heterophilic settings.
In this paper, we propose a novel approach, ClassContrast, grounded in Energy Landscape Theory from Chemical Physics, to overcome these limitations. ClassContrast combines spatial and contextual information, leveraging a physics-inspired energy landscape to model node embeddings that are both discriminative and robust across homophilic and heterophilic settings. Our approach introduces contrast-based homophily matrices to enhance the understanding of class interactions and tendencies. Through extensive experiments, we demonstrate that ClassContrast outperforms traditional GNNs in node classification and link prediction tasks, proving its effectiveness and versatility in diverse real-world scenarios.
Codehttps://anonymous.4open.science/r/CC-77A6/
@@ 1 Introduction
Over the past decade, GNNs have made a breakthrough in graph machine learning by adapting several deep learning models, originally developed for domains such as computer vision and natural language processing, to the graph setting (Section 2) . By using different approaches and architectures, GNNs produce powerful node embeddings by integrating neighborhood information with domain features. While recent GNN models consistently outperform the state-of-the-art (SOTA) results, in the seminal papers (Xu et al. 2019; Morris et al. 2019), the authors showed that the expressive power of message passing GNNs (MP-GNNs) is about the same as a well-known former method in graph representation learning, i.e. the Weisfeiler-Lehman algorithm (Shervashidze et al. 2011). Furthermore, recent studies show that GNNs suffer from significant challenges like oversquashing (Topping et al. 2021), oversmoothing (Oono and Suzuki 2020), and underreaching (Barceló et al. 2020).
@@
Another major issue with MPGNNs is their reliance on the (useful) homophily assumption, where edges typically connect nodes sharing similar labels and node features networks (Sen et al. 2008). However, many real-world scenarios exhibit heterophilic behavior, such as in protein and web networks (Pei et al. 2019) where conventional GNNs (Hamilton et al. 2017; Veličković et al. 2018a) may experience significant performance degradation, sometimes even underperforming compared to a multilayer perceptron (MLP) (Zhu et al. 2020; Luan et al. 2022). Therefore, there is a pressing need to develop models that function effectively in both homophilic and heterophilic settings.
We hypothesize that the limitations in current methods arise from the ineffective utilization of two key types of information in graphs: spatial information, which captures the local neighborhood’s class distribution for each node, and contextual information, which assesses each node’s similarity or dissimilarity to class representatives within a latent space, incorporating information from relevant but distant nodes. To address this issue, we leverage Energy Landscape Theory (Wales 2002; Wales et al. 1998) from Chemical Physics, a framework used to describe and analyze the behavior of complex systems, such as protein folding (Plotkin and Onuchic 2002). This theory visualizes the potential energy of a system as a multidimensional landscape, with valleys and peaks representing states of different energies.
We apply this theory to model a node’s position within a network, aiming to understand how nodes form connections based on specific contextual influences encoded as node attributes. In ClassContrast, integrating spatial and contextual information is similar to navigating an energy landscape where each node’s embedding seeks a low-energy state that is both highly discriminative and informative. In a homophilic setting, this state ideally balances similarity to nodes within the same class (lowering energy) and dissimilarity to nodes in different classes (avoiding overly general local minima). In a heterophilic setting, the model uses both types of information to address the complexities of non-conforming node relationships, where nodes that are dissimilar may still be connected. The energy landscape model helps identify and strengthen valuable connections that cross traditional boundaries, enhancing the model’s classification accuracy in environments with non-standard “like-likes-like” patterns. This dual capability allows ClassContrast to generate robust embeddings that perform well across varied graph structures, regardless of network homophily or heterophily.
Overall, ClassContrast effectively integrates spatial and contextual information to produce robust node embeddings, demonstrating impressive performance in node classification and link prediction tasks.
@@
    We propose a physics-inspired approach, leveraging Energy Landscape Theory, to construct class-aware node representations.
@@
    We introduce contrast-based homophily matrices that provide detailed insights into interactions and homophily tendencies among classes.
@@
    Our ClassContrast framework combines spatial and contextual information, achieving highly competitive results with state-of-the-art GNNs in both node classification and link prediction tasks.
@@
    Our experiments show that ClassContrast can be smoothly integrated with GNN models, boosting performance across both homophilic and heterophilic settings, proving its versatility across various real-world applications.
@@ 2 Related Work
Over the past decade, GNNs have dominated graph representation learning, especially excelling in node classification tasks (Xiao et al. 2022). After the success of Graph Convolutional Networks (GCN) (Kipf and Welling 2017), using a neighborhood aggregation strategy to perform convolution operations on graph, subsequent efforts focused on modifying, approximating, and extending the GCN approach, e.g., involving attention mechanisms (GAT) (Veličković et al. 2018a), sampling and aggregating information from a node’s local neighborhood (GraphSAGE) (Hamilton et al. 2017). However, a notable scalability challenge arose due to its dependence on knowing the full graph Laplacian during training. To address this limitation, numerous works emerged to enhance node classification based on GCN, such as methods , importance sampling node (FastGCN) (Chen, Ma, and Xiao 2018), adaptive layer-wise sampling with skip connections (ASGCN) (Huang et al. 2018), adapting deep layers to GCN architectures (deepGCN) (Li et al. 2019), incorporating node-feature convolutional layers (NFC-GCN) (Zhang et al. 2022).
@@
Two fundamental shortcomings of MPGNNs are the loss of structural information within node neighborhoods and the difficulty in capturing long-range dependencies. To address these inherent issues, numerous studies conducted in the past few years have introduced various enhancements. These improvements encompass diverse aggregation schemes, such as skip connections(Chen et al. 2020), geometric methods(Pei et al. 2019), aiming to mitigate the risk of losing crucial information. Additionally, advancements like implicit hidden layers (Geng et al. 2021) and multiscale modeling (Liu et al. 2022) have been explored to augment the GNN’s capabilities.
Another major issue with MPGNNs is that they operate under the homophily assumption and exhibit poor performance in heterophilic networks (Zhu et al. 2020; Luan et al. 2023). Hence, in the past few years, various efforts have been undertaken to develop GNNs that function effectively in heterophilic settings (Pei et al. 2019; Lim et al. 2021a; Luan et al. 2022; Zheng et al. 2022; Xu et al. 2023).
However, to the best of our knowledge, no previous work has developed a systematic approach to analyze class-wise relations.
3 Energy Landscape Theory and Its Application to Node Representation
Energy Landscape Theory originates from the field of chemical physics (Wales 2002), where it is utilized to describe the potential energy of a molecular system as a multidimensional landscape. This landscape is characterized by a series of valleys and peaks, each representing different states of energy. Valleys correspond to stable states (or low-energy configurations) where a system tends to reside under normal conditions, while peaks represent high-energy states that are less likely to be occupied.
In the context of graph neural networks, we apply the principles of Energy Landscape Theory to the problem of node representation learning. Here, we envisage nodes as entities navigating through an ’energy landscape’ constructed from the network structure and node attributes (see Appendix E for connections to the Structural Balance Theory (Heider 1946) on social networks). This landscape is shaped by the distribution of node features and the topological arrangement of the graph, wherein each node seeks a position corresponding to a state of minimal energy. For example, a fundamental result in social networks states that nodes aim to close friendship triangles (Granovetter 1973) to achieve a lower state (and foster trust). The absence of closed triangles indicates higher energy states, which may signify weaker relationships or even potential animosity. Such relationships are studied in signed network analysis using balance and structural theories (Leskovec, Huttenlocher, and Kleinberg 2010). We analogously define energy states by the optimal embedding that reflects accurate class membership and distinctiveness from nodes of different classes, which we measure with a class-aware definition of homophily (see Section 4.4). This approach allows us to conceptualize the embedding process as a dynamic system where nodes move towards lower-energy configurations that represent stable, informative, and discriminative embeddings in the context of the given graph.
Why is Energy Landscape Approach Useful?
We posit that a graph exhibits not merely a singular energy peak or valley, but multiple such features shaped by group formation dynamics. Consequently, nodes may display either homophilous or heterophilous behaviors depending on their group affiliations within a single graph. This variability implies that node representation should not uniformly assume proximity of neighbors as a reliable descriptor of a node’s characteristics. For example, Zhu et al. report that “in the Chameleon dataset, intra-class similarity is generally higher than inter-class similarity so MLP works better than GCN” (Zhu et al. 2020). Therefore, message-passing-based schemes should be employed selectively, based primarily on the presence of demonstrable homophily within groups. In the absence of such homophily, the system should learn to model heterophilic behaviors effectively. To this end, we claim that our framework, ClassContrast, is capable of simultaneously modeling both homophilous and heterophilous behaviors, adapting its approach as necessary for different groups.
We begin by detailing ClassContrast, focusing on two primary objectives: the discovery of spatial and contextual information. The first objective is to extract information from the node’s neighborhood (spatial), while the second is to derive insights from the node’s domain attributes (contextual). Our framework is visualized in Figure 1.
4 ClassContrast Embeddings
@@
Figure 1: ClassContrast framework. Spatial and contextual embeddings are incorporated in class-aware contrast-based learning.
@@
In the following discussion, we adhere to the notation $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{W},\mathcal{X})$ where $\mathcal{V}=\{v_{i}\}_{i=1}^{m}$ represents the vertices (nodes), =$\mathcal{E}=\{e_{ij}\}$ represents edges, $\mathcal{W}=\{\omega_{ij}\subset\mathbb{R}^{+}\}$ represents edge weights, and $\mathcal{X}=\{X_{i}\}_{i=1}^{m}\subset\mathbb{R}^{n}$ represents node features. If no node features are provided, $\mathcal{X}=\emptyset$. Consider a graph where nodes are categorized into $N$ classes. For each node $u$, we aim to construct an embedding $\vec{\gamma}(u)$ of dimension $q\times N$, where $q$ is determined by specifics of the graph such as its directedness, weighted nature, and the format of domain features.
@@
Specifically, we generate spatial embeddings $\{\vec{\alpha}_{1}(u),\vec{\alpha}_{2}(u),\dots\}$, derived from local neighborhood information, and contextual embeddings $\{\vec{\beta}_{1}(u),\vec{\beta}_{2}(u),\dots\}$, derived from node properties specific to the domain. Each embedding, $\vec{\alpha}_{i}(u)$ and $\vec{\beta}_{j}(u)$, is $N$-dimensional, with one entry corresponding to each class (ClassContrast coordinates). These embeddings are then concatenated to form the final embedding $\vec{\gamma}(u)$ of size $q\times N$. For simplicity, we focus on applications in an inductive setting. In Section C.2, we discuss adaptations of our methods to accommodate transductive settings.
4.1 Spatial Node Embedding
@@
Assume that there are $N$ classes of nodes, represented as $\mathcal{C}_{1},\mathcal{C}_{2},\dots,\mathcal{C}_{N}$. Let $\mathcal{V}_{\text{train}}$ be the set of nodes with known labels in the training dataset.
[…]
@@
4.2 Contextual Node Embedding
Inspired by Energy Landscapes, contextual node embeddings utilize the attribute space to measure the distance of a node to known classes. In social networks, node attributes might include user-specific details (Mislove et al. 2007), while in citation networks, they often involve the presence or absence of certain keywords (Tang et al. 2008). These attributes are crucial for predicting node behavior as they provide intrinsic information about the nodes themselves. Hence, we develop a general and versatile approach to effectively utilize this information in node classification task.
We use the term contextual node embeddings to highlight the importance of context within the graph structure, beyond isolated node attributes. Unlike attribute embeddings, which embed node characteristics directly, contextual node embeddings integrate the node’s attributes and its relational and structural position in the graph. This is crucial in scenarios where class behavior and node features interact, such as an article’s topic influencing its citation links.
A real-life context example.
Consider the CORA dataset where publications that are frequently cited by other papers, especially by those within the same class (topic), may be considered authoritative in their field. An embedding that captures this citation context can differentiate between a highly influential paper and one that is similarly cited (by papers from other classes) but less authoritative.
In contextual embeddings we identify two cases on attribute availability and quality. First, attributes may already be available in high quality vectors, such as word embeddings (e.g., word2vec (Mikolov et al. 2013a, b)). We directly integrate these vectors without additional processing, as detailed in Appendix D. When node attributes are available but unprocessed, we define an attribute vector $\mathcal{X}(u)$ for each node $u$, which maps nodes to an $n$-dimensional real vector space. For each class $\mathcal{C}_{j}$, we identify a cluster of points $\mathcal{Z}_{j}=\{\mathcal{X}(u)\mid u\in\mathcal{C}_{j}\}$ and establish a representative landmark $\xi_{j}$ in $\mathbb{R}^{n}$ for this cluster.
@@ 4.2 Contextual Node Embedding
Given a node class $\mathcal{C}_{j}$ and its corresponding cluster of node embeddings $\mathcal{Z}_{j}$, the landmark $\xi_{j}$ is computed as the centroid of the points in $\mathcal{Z}_{j}$, normalized by the number of nodes in $\mathcal{C}_{j}$: $\xi_{j}=\frac{1}{|\mathcal{C}_{j}|}\sum_{u\in\mathcal{C}_{j}}\mathcal{X}(u).$
@@
Given a node $u$ and a set of class landmarks $\{\xi_{j}\}\subseteq\mathbb{R}^{n}$, one for each class $\mathcal{C}_{j}$, the contextual embedding of node $u$ is defined by a vector of distances from the node’s embedding to each class landmark: $\vec{\beta}(u)=[d_{1},d_{2},d_{3},\dots,d_{N}]$ where $d_{j}=\mathbf{d}(\mathcal{X}(u),\xi_{j})$ for each class $\mathcal{C}_{j}$ (see Figure 2(b)). This vector $\vec{\beta}(u)$ encapsulates the node’s position relative to each class within the attribute space.
Landmark Sets.
To obtain information about the domain attributes of nodes for each class, multiple landmarks can be defined. For instance, while the first landmark $\xi^{1}_{j}\in\mathbb{R}^{n}$ represents the center of each cluster $\mathcal{Z}_{j}$, the second landmark $\xi^{2}_{j}\in\mathbb{R}^{n}$ may capture the most frequent attributes of $\mathcal{Z}_{j}$ (see Appendix D). For each type of landmark, a corresponding distance or similarity measure $\mathbf{d}^{k}(.,.)$ is defined, such as Euclidean or cosine distance for real-valued attribute vectors, or Jaccard similarity for categorical attributes (Niwattanakul et al. 2013). Subsequently, another $N$-dimensional vector $\vec{\beta}_{k}(u)=[d_{k1},d_{k2},d_{k3},\dots,d_{kN}]$ is obtained, where $d_{kj}=\mathbf{d}^{k}(\mathcal{X}(u),\xi^{k}_{j})$ for each class $\mathcal{C}_{j}$. Landmark set extensions are explained in Appendix D.
4.3 ClassContrast Embedding
We may expand spatial neighborhoods, and extend landmark sets arbitrarily. However, for exposition purposes, we will define ClassContrast embeddings over a directed graph where each node’s neighborhood is considered up to two hops and each class $\mathcal{C}_{j}$ has two landmarks. We define the ClassContrast Embedding of a node $u$ by concatenating spatial and contextual embeddings. Specifically, we consider:

### v2

# SCNode: Spatial and Contextual Coordinates for Graph Representation Learning

@@ Abstract
Effective node representation lies at the heart of Graph Neural Networks (GNNs), as it directly impacts their ability to perform downstream tasks such as node classification and link prediction. Most existing GNNs, particularly message passing graph neural networks, rely on neighborhood aggregation to iteratively compute node embeddings. While powerful, this paradigm suffers from well-known limitations of oversquashing, oversmoothing, and underreaching that degrade representation quality. More critically, MPGNNs often assume homophily, where connected nodes share similar features or labels, leading to poor generalization in heterophilic graphs where this assumption breaks down.
To address these challenges, we propose SCNode, a Spatial-Contextual Node Embedding framework designed to perform consistently well in both homophilic and heterophilic settings. SCNode integrates spatial and contextual information, yielding node embeddings that are not only more discriminative but also structurally aware. Our approach introduces new homophily matrices for understanding class interactions and tendencies. Extensive experiments on benchmark datasets show that SCNode achieves superior performance over conventional GNN models, demonstrating its robustness and adaptability in diverse graph structures.
@@ 1 Introduction
Over the past decade, GNNs have made a breakthrough in graph machine learning by adapting several deep learning models originally developed for domains such as computer vision and natural language processing to the graph setting. By using different approaches and architectures, GNNs produce powerful node embeddings by integrating neighborhood information with domain features. While recent GNN models consistently outperform the state-of-the-art (SOTA) results, in the seminal papers (Xu et al., 2019; Morris et al., 2019), the authors showed that the expressive power of message passing GNNs is about the same as a well-known former method in graph representation learning, i.e., the Weisfeiler-Lehman algorithm (Shervashidze et al., 2011). Furthermore, recent studies show that GNNs suffer from significant challenges such as oversquashing (Topping et al., 2021), oversmoothing (Oono & Suzuki, 2020), and underreaching (Barceló et al., 2020).
@@
Another major issue with MPGNNs is their reliance on the (useful) homophily assumption, where edges typically connect nodes sharing similar labels and node attributes (Sen et al., 2008). However, many real-world scenarios exhibit heterophilic behavior, such as in protein and web networks (Pei et al., 2019), where conventional GNNs (Hamilton et al., 2017; Veličković et al., 2018a) may experience significant performance degradation, sometimes even underperforming compared to a multilayer perceptron (MLP) (Zhu et al., 2020; Luan et al., 2022; 2024). Therefore, there is a pressing need to develop models that function effectively in both homophilic and heterophilic settings.
We attribute current GNN shortcomings to their inability to capture two complementary sources of information in graphs: spatial information, which reflects the class distribution in a node’s immediate neighborhood, and contextual information, which measures a node’s relation to class representatives across the entire graph. To address this, we position each node relative to all classes. In the attribute space, we introduce class landmarks, reference embeddings that act as beacons for each class, and record the distances from each node’s feature vector to these landmarks as its contextual coordinates. In the graph topology, we encode spatial coordinates by computing the distribution of neighboring node labels within each node’s local subgraph.
Our landmark scheme establishes an implicit coordinate system: by triangulating distances to class landmarks, we locate each node in attribute space much like a GPS. These relative coordinates capture global patterns that standard message-passing GNNs often miss due to underreaching, i.e., the inability to propagate sufficient information from distant but relevant nodes within a limited number of layers. Spatial embeddings encode local label distributions in multi-hop neighborhoods, enriching high-frequency structural signals, while contextual embeddings inject low-frequency global components through landmark-based coordinates. From a spectral perspective, concatenating these embeddings expands the effective receptive field by combining complementary frequency bands. This joint representation mitigates the 1-WL locality bottleneck and enhances expressivity by unifying local structural distributions with long-range semantic context.
By leveraging relative positioning, we generate feature vectors that seamlessly blend spatial and contextual cues. The resulting SCNode model achieves state-of-the-art performance on node classification and link prediction across both homophilic and heterophilic benchmarks. Furthermore, SCNode embeddings can be dropped into any GNN as plug-and-play components, remedying underreaching and consistently boosting performance in diverse graph settings.
@@
    We propose a novel approach to capture the relative positioning of nodes with respect to node classes, incorporating both spatial and contextual perspectives.
@@
    We introduce class-aware homophily matrices, providing detailed insights into homophily tendencies and enabling a deeper understanding of class interactions in the graph.
@@
    Our SCNode model effectively integrates spatial and contextual information, delivering state-of-the-art performance in both node classification and link prediction tasks.
@@
    We demonstrate that SCNode vectors can be easily integrated into any GNN model as plug-and-play components, significantly boosting performance in both homophilic and heterophilic settings, showcasing their versatility across diverse real-world applications.
@@ 2 Related Work
Over the past decade, GNNs have dominated graph representation learning, especially excelling in node classification tasks (Xiao et al., 2022). After the success of Graph Convolutional Networks (GCN) (Kipf & Welling, 2017), using a neighborhood aggregation strategy to perform convolution operations on graph, subsequent efforts focused on modifying, approximating, and extending the GCN approach, e.g., involving attention mechanisms (GAT) (Veličković et al., 2018a), sampling and aggregating information from a node’s local neighborhood (GraphSAGE) (Hamilton et al., 2017). However, a notable scalability challenge arose due to its dependence on knowing the full graph Laplacian during training. To address this limitation, numerous works emerged to enhance node classification based on GCN, such as methods, importance sampling node (FastGCN) (Chen et al., 2018), adaptive layer-wise sampling with skip connections (ASGCN) (Huang et al., 2018), adapting deep layers to GCN architectures (deepGCN) (Li et al., 2019), incorporating node-feature convolutional layers (NFC-GCN) (Zhang et al., 2022).
@@
Two fundamental shortcomings of MPGNNs are the loss of structural information within node neighborhoods and the difficulty in capturing long-range dependencies (Khemani et al., 2024; Corso et al., 2024). To address these inherent issues, numerous studies conducted in the past few years have introduced various enhancements, including diverse aggregation schemes, such as skip connections (Chen et al., 2020), geometric methods (Pei et al., 2019), aiming to mitigate the risk of losing crucial information. Additionally, advancements like implicit hidden layers (Geng et al., 2021) and multiscale modeling (Liu et al., 2022) have been explored to augment the GNN’s capabilities.
Another major issue with MPGNNs is their reliance on the homophily assumption, leading to poor performance in heterophilic networks (Zhu et al., 2020; Luan et al., 2024). Recent efforts have focused on designing GNNs that perform well in heterophilic settings (Lim et al., 2021a; Luan et al., 2022; Zheng et al., 2022; Xu et al., 2023; Zhao et al., 2024). While prior works analyze class-wise relations via neighborhood label distributions (Ma et al., 2021; Zhu et al., 2023) or class-compatibility matrices (Zhu et al., 2021), SCNode introduces a coordinate-based formulation that combines local structural distributions with global class-aware landmark distances, providing backbone-agnostic feature augmentations that capture long-range class relations.
In addition, graph kernels, such as Weisfeiler-Lehman (Shervashidze et al., 2011), shortest-path Borgwardt & Kriegel (2005), and graphlet kernels (Shervashidze et al., 2009), characterize similarity through handcrafted structural mappings (Vishwanathan et al., 2010). SCNode is related in spirit, but it differs in two important ways. It combines local structural distributions with global class-aware relational signals, providing richer contextual information than fixed kernel similarities. Moreover, SCNode is computationally more efficient, avoiding the quadratic complexity of kernel evaluations by generating embeddings in linear time that can be directly used within GNNs or graph transformers.
3 SCNode Framework
In the following discussion, we use the notation $\mathcal{G}=(\mathcal{V},\mathbb{E},\mathcal{W},\mathcal{X})$ where $\mathcal{V}=\{v_{i}\}_{i=1}^{m}$ represents
@@ 3 SCNode Framework
Figure 1: SCNode framework. Spatial and contextual embeddings are incorporated in class-aware contrast-based learning.
@@
the vertices (nodes), $\mathbb{E}=\{e_{ij}\}$ represents edges, $\mathcal{W}=\{\omega_{ij}\subset\mathbb{R}^{+}\}$ represents edge weights, and $\mathcal{X}=\{\mathcal{X}_{i}\}_{i=1}^{m}\subset\mathbb{R}^{n}$ represents node features. If no node features are provided, $\mathcal{X}=\emptyset$. Consider a graph where nodes are categorized into $N$ classes. For each node $u$, we aim to construct an embedding $\vec{\gamma}(u)$ of dimension $q\times N$, where $q$ is determined by specifics of the graph such as its directedness, weighted nature, and the format of domain features.
@@
Specifically, we generate spatial embeddings $\{\vec{\alpha}_{1}(u),\vec{\alpha}_{2}(u),\dots\}$, derived from local neighborhood information, and contextual embeddings $\{\vec{\beta}_{1}(u),\vec{\beta}_{2}(u),\dots\}$, derived from node properties specific to the domain. Each embedding, $\vec{\alpha}_{i}(u)$ and $\vec{\beta}_{j}(u)$, is $N$-dimensional, with one entry corresponding to each class (SCNode coordinates). These embeddings are then concatenated to form the final embedding $\vec{\gamma}(u)$ of size $q\times N$. For simplicity, we focus on applications in an inductive setting. In Section˜B.2, we discuss adaptations of our methods to accommodate transductive settings.
3.1 Spatial Node Embeddings
@@
Assume that there are $N$ classes of nodes, represented as $\mathcal{C}_{1},\mathcal{C}_{2},\dots,\mathcal{C}_{N}$. Let $\mathcal{V}_{\text{train}}$ be the set of nodes with known labels in the training dataset. The feature vector $\vec{\alpha}_{k}(u)$ for node $u$ is initialized as the distribution of class occurrences: $\vec{\alpha}_{k}(u)=[a_{k1},a_{k2},\dots,a_{kN}]$ where $a_{kj}$ counts how many neighbors of $u$ belong to class $\mathcal{C}_{j}$ within the k-hop neighborhood intersecting with $\mathcal{V}_{\text{train}}$: $a_{kj}=|\{v\in\mathcal{C}_{j}\mid v\in\mathcal{N}_{k}(u)\cap\mathcal{V}_{\text{train}}\}|$. Consider the toy example of Figure 2, where the (k=1)-hop neighborhood of node $u$ contains node $z$ from class $1$ and node $y$ from class $2$. As a result, $\vec{\alpha}_{1}(u)=[1\ 1\ 0]$. We extend the spatial embeddings to directed and weighted graph cases by incorporating edge directions and weights in additional dimensions (see Appendix Section B.3) for the definitions.
3.2 Contextual Node Embeddings
In this section, we focus on extracting the most relevant information from the node attribute space by utilizing class-level reference points. Instead of relying solely on local neighborhood information, we assess each node’s position relative to a set of representative class embeddings, or class landmarks, defined in the attribute space. These landmarks serve as stable points that reflect the typical feature profiles of different classes.
To compute contextual embeddings, we measure the similarity or distance between a node’s attribute vector, $\mathcal{X}_{u}$, and these class landmarks. This approach enables the model to capture how closely a node aligns with each class, regardless of its immediate neighbors.
We use the term contextual node embeddings to emphasize the integration of node attributes with relational and structural information from the graph. Unlike traditional attribute embeddings, which focus solely on
@@ 3.2 Contextual Node Embeddings
embedding node characteristics, contextual embeddings account for both the intrinsic properties of a node and its position within the graph structure. This approach is especially important in scenarios where node features and class behavior are intertwined, such as when the topic of an article influences its citation links. By combining attribute and relational context, our embeddings provide a richer representation of the graph, enabling improved performance in classification tasks.
In the following, we define class representatives (landmarks) in the node attribute space for each class. Let $\{\mathcal{C}_{j}\}_{j=1}^{N}$ denote the set of node classes. For each class $\mathcal{C}_{j}$, we identify the cluster of points $\mathcal{Z}_{j}=\{\mathbf{X}(u)\mid u\in\mathcal{C}_{j}\}$ within the node attribute space $\mathbb{R}^{n}$, and establish a representative landmark $\xi_{j}\in\mathbb{R}^{n}$ for this cluster.
@@
Given a node class $\mathcal{C}_{j}$ and corresponding cluster of node embeddings $\mathcal{Z}_{j}$, the landmark $\xi_{j}$ is computed as the centroid of the points in $\mathcal{Z}_{j}$, normalized by the number of nodes in $\mathcal{C}_{j}$:
$\xi_{j}=\dfrac{1}{|\mathcal{C}_{j}|}\sum_{u\in\mathcal{C}_{j}}\mathcal{X}(u)$
@@
Given a node $u$ and a set of class landmarks $\{\xi_{j}\}\subseteq\mathbb{R}^{n}$, one for each class $\mathcal{C}_{j}$, the contextual embedding of node $u$ is defined by a vector of distances from the node’s embedding to each class landmark:
$\vec{\beta}(u)=[d_{1},d_{2},d_{3},\dots,d_{N}]$
where $d_{j}=\mathbf{d}(\mathcal{X}(u),\xi_{j})$ for each class $\mathcal{C}_{j}$ (see Figure 2). This vector $\vec{\beta}(u)$ encapsulates the node’s position relative to each class within the attribute space.
Class Landmarks.
To extract more detailed information about the domain attributes of nodes for each class, multiple landmarks can be defined. In many cases, the node attribute space lacks an inherent metric, and using different metrics can result in diverse landmarks and contextual embeddings. For example, the first landmark $\xi^{1}_{j}\in\mathbb{R}^{n}$ may represent the center of the cluster $\mathcal{Z}_{j}$ using the Euclidean metric, while a second landmark $\xi^{2}_{j}\in\mathbb{R}^{n}$ for the same class can be based on Jaccard similarity, capturing the most frequent attributes within $\mathcal{Z}_{j}$ (see Appendix˜C).
For each type of landmark, a corresponding distance or similarity measure $\mathbf{d}^{k}(.,.)$ is defined, such as Euclidean or cosine distance for real-valued attribute vectors, or Jaccard similarity for categorical attributes (Niwattanakul et al., 2013).
By leveraging different metrics and their corresponding landmarks, we compute an $N$-dimensional vector
$\vec{\beta}_{k}(u)=[d_{k1},d_{k2},d_{k3},\dots,d_{kN}]$
where $d_{kj}=\mathbf{d}^{k}(\mathbf{X}(u),\xi^{k}_{j})$ represents the distance (or similarity) between node embeddings $\mathbf{X}(u)$ and the landmark $\xi^{k}_{j}$ of class $\mathcal{C}_{j}$.
3.3 SCNode Embedding
We may expand spatial neighborhoods and extend landmark sets arbitrarily. However, for exposition purposes, we will define SCNode embeddings over a directed graph where each node’s neighborhood is considered up to two hops and each class $\mathcal{C}_{j}$ has two landmarks. We define the SCNode Embedding of a node $u$ by concatenating spatial and contextual embeddings. Specifically, we consider: